OpenStudy (anonymous):

find the ordered pair solution to the following system of equations: 3x -2y = -1 2x + y = y

OpenStudy (anonymous):

Hmm... Unless there's a typo... The second equation 2x + y = y gives us 2x = 0, so x must be 0. From that, we can plug x = 0 into the first equation, and see that -2y = -1. So y = 1/2. So (0, 1/2) is our ordered pair solution.

OpenStudy (anonymous):

that was suppose to have been 2x + y = 4

OpenStudy (anonymous):

im sorry. can u please help me with that

OpenStudy (anonymous):

what would the answer be with the y being a 4 instead of y

OpenStudy (anonymous):

when 2x+y = 4 then 7x = 7 (after multiplying second by 2 and add to first) x =1 and y = 2 so, (1,2) is a solution......